# Advancing high-fidelity 3D and Texture Generation with 2.5D latents

### v1

@@ Abstract.
text and image to 3D generation, text and image to texture generation
journal: TOG
Figure 1. Examples of “X” to 3D and “X” to texture generation. In this paper, we propose a new approach to bridge 3D generation with 2D diffusion priors, i.e., the generation of 2.5D latent. By leveraging the advantage of 3D-like representation and the prior of pretrained 2D diffusion models to the best extent, we achieve not only 3D generation with high-fidelity, but also excels in geometry-conditioned texture generation with image or text conditions. Please zoom in for details.
00footnotetext: * Equal contribution
@@ 1. Introduction
For example, we observe that most state-of-the-art 3D generation models (Li et al., 2025; Team, 2025; Zhang et al., 2024; Li et al., 2024) adopt a two-stage 3D generation pipeline, consisting of a geometry generation stage followed by a texture generation stage. During the texture generation stage, one must exert additional efforts (Huang et al., 2024; Xiong et al., 2024; Zheng et al., 2024) to ensure that the generated texture aligns with the pre-generated geometry.
A recent approach, TRELLIS (Xiang et al., 2024), seeks to integrate geometry and texture generation through a 3D structured latent (SLat) generation pipeline. Although TRELLIS is capable of producing highly aligned 3D assets with refined structures, we find that the texture quality of its outputs is often inferior to that of multi-view-based 3D generative models (Lin et al., 2025; Huang et al., 2024; Wu et al., 2024), which utilize 2D generative priors.
@@
These multiview images in necessary modalities forms the 2.5D representation of the object. With the coordinate components of the 2.5D representation, we can easily project the multiview features into the structured 3D voxel space, which enables 3D native asset reconstruction. Compared with the native 3D representations such as voxel grid and triplane, the 2.5D representation is essentially images, and thus much easier to process. Hence, we can generate such 2.5D latents by adapting 2D multiview generative models as previous works (Lin et al., 2025; Wu et al., 2024; Long et al., 2023).
@@
Figure 2. The multiview 2.5D representation. In this paper, we curate a 2.5D dataset by rendering the multiview RGB, normal and coordinate maps from 3D assets. With the dataset, we encode the image of each modalities into 2D latents with 2D VAE, and project the latents into 3D voxel space with multiview averaging. Then, we train our 3D refiner and decoder to reconstruct the 3D assets (3D gaussian splats or mesh) from the aggregated structured latents.
2. Related Works
2.1. 2D Diffusion Models
@@ 2. Related Works
Diffusion models (Sohl-Dickstein et al., 2015; Ho et al., 2020) define a Markov process that gradually transforms data into noise through a forward diffusion process and then learns to reverse this transformation to generate data from noise. They have demonstrated remarkable performance in image (Rombach et al., 2022; He et al., 2024b; Ramesh et al., 2021; Saharia et al., 2022; He et al., 2024a) and video generation (Ho et al., 2022a; Blattmann et al., 2023a, b; Ho et al., 2022b). However, conventional diffusion trajectories often exhibit significant curvature (Karras et al., 2022; Liu et al., 2022), which increases the number of required solver steps and amplifies error accumulation. To address this, Rectified Flow (Liu et al., 2022; Albergo and Vanden-Eijnden, 2022; Lipman et al., 2023) has been proposed, enforcing a straight-line trajectory between noise and data. Flow-based diffusion models have shown strong empirical performance and high efficiency, as demonstrated by recent state-of-the-art systems such as Stable Diffusion V3(Esser et al., 2024) and Flux.1 (BlackForestLabs, 2024). In this work, we adopt Flux.1 as the backbone diffusion model for 2.5D generation, leveraging its rectified flow design to enable high-quality and efficient image synthesis, which serves as a strong prior in our 3D reconstruction pipeline.
2.2. 3D Generative Models
2.2.1. Multiview Diffusion Models
@@
Multiview diffusion models (Zheng et al., 2024; Lu et al., 2024; Miao et al., 2024; Qin et al., 2025; Chen et al., 2024; Lin et al., 2025; Shi et al., 2023; Xu et al., 2024) have focused on synthesizing geometrically consistent object views from varying camera poses, conditioned on image or text prompts. Early efforts in multiview generation models, such as MVDream (Shi et al., 2023), have successfully adapted pre-trained text-to-image diffusion models to generate object-centric multiview images. To enhance the geometric fidelity of the generated 3D content, methods such as Direct2.5 (Lu et al., 2024) and Kiss3DGen (Lin et al., 2025) incorporate geometry-aware designs and normal-guided supervision, producing multiview normal maps that facilitate high-quality 3D asset reconstruction. Other approaches leverage explicit 3D representations to improve realism and structural integrity.  Miao et al. (2024); Qin et al. (2025) distill multiview diffusion outputs into compact 3D generators using Gaussian splatting (Kerbl et al., 2023), enabling efficient single view 3D reconstruction with enhanced visual quality. In addition, auxiliary modules have been proposed to improve cross-view consistency and 3D reasoning.  Chen et al. (2024) introduces a modular architecture called 3D-Adapter that injects 3D priors into diffusion pipelines, while  Zheng et al. (2024) constructs a volumetric feature representation by aggregating multiview image features, leading to more accurate geometry reconstruction.
2.2.2. 3D Diffusion Models
@@
Xiang et al. (2024) introduces a unified structured latent space for 3D content (SLAT), supporting diverse representations like meshes, Gaussians, and NeRF.  Zhang et al. (2024) uses a DiT-based latent diffusion model to generate neural surface representations and materials. It combines a multi-resolution VAE with PBR-aware training and supports conditional generation from text or images.  Li et al. (2024) performs native 3D diffusion directly in volumetric space to generate coarse mesh shapes. It integrates a normal-guided surface refiner that enhances geometry.  Liu et al. (2023) directly applies score-based diffusion to mesh vertex coordinates and face connectivity. It models 3D meshes in a native representation, enabling faithful and topology-preserving mesh generation without intermediate voxelization or rendering supervision.
Figure 3. Our proposed framework. In this paper, we propose a unified framework for (b) text or image to 3D generation and (c) geometry-conditioned texture generation. To provide the (a) hybrid image-text condition for 3D or texture generation, we adopt the off-the-shelf image encoder SigLIP (Zhai et al., 2023) and Flux.1-Redux (BlackForestLabs, 2024) image embedder with the T5 and CLIP text encoder. During the training, we randomly dropout the image or text condition to maintain the model’s ability to generate coherent content with ”X“ conditions and achieve better model performance, please refer to our experiment section for more details.
3. Method
@@
3.1. Bridging 3D and 2D with 2.5D Representation
In our framework, we utilize an image-like 2.5D representation (Fig. 2, stage I) for object encoding and decoding. Unlike 3D representations, the 2.5D representations are essentially images, however, it combines the multiview RGB, normal and coordinate images to represent the volume, surface and texture information that are sufficient for high-quality 3D objects reconstruction. Although some previous works (Lin et al., 2025; Wu et al., 2024; Long et al., 2023; Wang et al., 2024) have explored the potential of similar representations that composed of rgb with either coordinate or normal map, the absence of either normal or coordinate map in their framework could introduce extra difficulties in the 3D reconstruction stage.
From 2.5D to 3D. While most previous works adapted sparse-view reconstruction approaches to reconstruct the object from either rgb, normal or coordinate information, their result often limited to specific 3D representation, and sometimes low-quality in the unseen areas of the multi-view results. Inspired by the 3D structured latent representation proposed by Trellis (Xiang et al., 2024), we also create a structure latent representation, which can then be decoded to various 3D representations in high-quality. For pipeline simplicity, we choose the VAE of the 2D diffusion model (i.e. Flux.1-dev) to encode the 2.5D images into 2D latents. Then, we project the 2D latents with $C$ channels and $N$ views $x_{2d}\in R^{N\times 3\times C\times H\times W}$ into a 3D latent voxel grid $x_{3d}\in R^{3C\times X\times Y\times Z}$ accordingly. As shown in stage II of Fig. 2, the projection of the 2D latent introduces a voxel-like 3D latent representation that contains the color, surface and coordinate information. Following (Xiang et al., 2024), we aim to introduce a sparse 3D decoder to reconstruct the structured latent into explicit 3D representations such as 3D gaussian splats or mesh.
@@
(1) $\hat{x}_{occ}=\begin{cases}1,&\text{if}~{}x_{occ}+b_{occ}>0.5\\ 0.&\text{otherwise}\end{cases}$
Then, we train the decoder to convert the refined structure latent into target 3D representation such as 3D gaussian splatting or mesh. Following  (Xiang et al., 2024), we adopt the sparse transformer for decoder, and train the refiner unet and 3DGS decoder in an end-to-end manner from scratch with rendering loss only. For the mesh decoding, we fixed the pretrained refiner and train the mesh decoder only. Please refer to our appendix for more details.
3.2. Model Fine-tuning Via Mixture-of-LoRA
Inspired by Lin et al. (2025), we aims to finetune a pretrained 2D diffusion model, i.e. Flux.1-dev (BlackForestLabs, 2024), via Low-Rank Adaptation (LoRA) (Hu et al., 2022) for 2.5D image generation. As our 2.5D image comprise RGB, normal, and coordinate images, extending the 2D diffusion model to the joint generation of multi-modal images using LoRA adapters with limited data presents a challenge. Furthermore, we observe that simply finetune a single LoRA adapter with large rank (e.g. 384) do not achieves the best performance, both quantitatively and qualitatively (see Sec. 9).
Figure 4. The architecture of proposed modules. In order to fix the occluded areas in the structure latent aggregated from our 2.5D latent, we introduce the (a) 3D Residual UNet to refine the feature and occupancy field. Then, we apply the (b) sparse transformer decoder (Xiang et al., 2024) for 3DGS or mesh reconstruction. (c) For each MLP block, we introduce three separate LoRA adapters tailored for 2.5D latent generation. Specifically, while the general LoRA layer is shared by all modalities, we feed the normal and coordinate (coord.) features into the auxiliary LoRA layers for extra feature projection, then we merge the results of the general and normal or coord. LoRAs via a simple addition.
@@ 3.2. Model Fine-tuning Via Mixture-of-LoRA
(2) $h_{i}=\begin{cases}F(x_{i})+\hat{F}_{\text{general}}(x_{i}),&\text{if}~{}i=\text{``rgb''}\\ F(x_{i})+\hat{F}_{\text{general}}(x_{i})+\hat{F}_{i}(x_{i}).&\text{otherwise}\end{cases}$
@@
Joint training of text and image condition. Unlike previous works, we build a versatile framework for text and image conditioned generation of the 2.5D latent. As we chose the pretrained 2D diffusion model (i.e., Flux-dev.1) as the base model for 2.5D generation, we adopt the off-the-shelf text encoder (T5 and CLIP) and image encoder (SigLIP (Zhai et al., 2023) image encoder with Flux-Redux (BlackForestLabs, 2024) image embedder) to embed the text and image conditions respectively. During the model finetuning, we randomly drop the image or text condition, which we found beneficial to both text and image conditioned generation tasks. Please refer to our experiment section for more details.
4. Experiment
4.1. Training details
@@
4.2. Dataset
Due to the heterogeneous quality of the original Objaverse dataset, we initially excluded assets that lacked texture maps or exhibited low mesh fidelity. A comprehensive manual curation process was then performed to remove low-quality samples, including incomplete geometries, scanned planar surfaces, and large-scale scene compositions. Additionally, to address inconsistencies in object orientations, we manually annotated the canonical front-facing direction for each object. Then, we further apply a filtering following Trellis (Xiang et al., 2024) to remove the items with low aesthetic score.
@@ 4.2. Dataset
For data rendering, we used Blender with a fixed camera distance of 4.5 units and a field of view (FoV) of 30 degrees. Each object was rendered from six distinct views to construct our 2.5D representation: four views were captured at 90-degree intervals in azimuth with a fixed elevation of 5 degrees—starting from the annotated front-facing direction—and an additional top-down view was rendered from a 90-degree elevation aligned with the frontal axis. For each view, we rendered RGB images, surface normal maps, and coordinate maps at a resolution of 512×512 pixels. The RGB multi-view images were then used as inputs for generating 3D-aware textual annotations via GPT-4V and Florence2 (Xiao et al., 2023).
4.3. Evaluation
For evaluation, we conducted both quantitative and qualitative comparisons. Quantitatively, we used the Google Scanned Objects (GSO) dataset (Downs et al., 2022). For the text-to-3D and text-to-texture task, we randomly sampled 100 objects and generated four-view renderings with GPT-4V annotations as text prompts. For image-to-3D and image-to-texture, we sampled 200 objects and rendered a single front-facing view as input for evaluation.
@@ 4.3. Evaluation
4.3.1. Metrics
We employ a comprehensive set of quantitative metrics to evaluate both the visual fidelity and geometric accuracy of the generated 3D models. For the text-to-3D and text-to-texture task, where precise 3D ground truth is unavailable, we assess semantic alignment using the CLIP score (Hessel et al., 2021) by rendering four orthogonal views of each generated object. In addition, we use Q-Align (Wu et al., 2023), a large multi-modal evaluation model, to assess the quality and aesthetic consistency of the rendered outputs.
@@ 4.3.1. Metrics
4.4. Comparison with State-of-the-Art Methods
In this section, we compare our approach with several state-of-the-art methods across three tasks: text-to-3D generation, image-to-3D generation and texture generation. For text-to-3D generation, we compared our method with four representative baselines: the diffusion-based model CraftsMan (Li et al., 2024) and Trellis (Xiang et al., 2024), Hunyuan3D-1 (Team, 2024) which builds on a large reconstruction model (LRM), Unique3D (Wu et al., 2024), a two stage generation method. For image-to-3D generation, We compared our method against three baselines: the latent diffusion model 3DTopia (Hong et al., 2024), Hunyuan3D-1 (Team, 2024), and Trellis (Xiang et al., 2024). For texture generation, we compared our method against Hunyuan2.0 (Team, 2025), MV-Adapter (Huang et al., 2024) which generates a single image from text, and then expand it into six multi-view images in image-to-texture generation, and SyncMVD (Liu et al., 2024) which generates a central image from text and predicts synchronized multi-view images via a cross-view fusion module, TexGaussian (Xiong et al., 2024) using 3D Gaussians in text-to-texture generation.
Image-to-3D generation. As shown in Table 1, our method achieves competitive performance across both geometric and perceptual metrics. Notably, our results are comparable to Trellis, despite Trellis relying on significantly larger-scale training data and a much larger model. In contrast, our method is designed to be more data- and compute-efficient, yet achieves the highest PSNR and a competitive LPIPS score, ranking closely behind Trellis. Our SSIM score is also comparable to top-performing methods, indicating strong perceptual consistency and visual fidelity in the rendered views. These results suggest that our approach effectively reconstructs high-quality 3D shapes from single-view inputs, achieving a favorable trade-off between visual quality and computational cost. In addition to the quantitative results, Figure 5 presents a qualitative comparison under the image-to-3D setting. Compared to TripoSG (Li et al., 2025) and Trellis, our method produces significantly more faithful and detailed textures. For example, our results better preserve color contrast, facial features, and fine-grained patterns such as fur, eyes, and accessories. Furthermore, our textures remain consistent across multiple viewpoints, reflecting better integration between appearance and structure.

### v2

@@ Abstract
Figure 1: Examples of “X” to 3D and “X” to texture generation. In this paper, we propose a new approach to bridge 3D generation with 2D diffusion priors, i.e., the generation of 2.5D latent. By leveraging the advantage of 3D-like representation and the prior of pretrained 2D diffusion models to the best extent, we achieve not only 3D generation with high-fidelity, but also excels in geometry-conditioned texture generation with image or text conditions. Please zoom in for details.
footnotetext: * Equal contribution
@@ 1 Introduction
For example, we observe that most state-of-the-art 3D generation models [20, 37, 46, 19] adopt a two-stage 3D generation pipeline, consisting of a geometry generation stage followed by a texture generation stage. During the texture generation stage, one must exert additional efforts [16, 43, 47] to ensure that the generated texture aligns with the pre-generated geometry.
A recent approach, TRELLIS [41], seeks to integrate geometry and texture generation through a 3D structured latent (SLat) generation pipeline. Although TRELLIS is capable of producing highly aligned 3D assets with refined structures, we find that the texture quality of its outputs is often inferior to that of multi-view-based 3D generative models [21, 16, 40], which utilize 2D generative priors.
@@
These multiview images in necessary modalities forms the 2.5D representation of the object. With the coordinate components of the 2.5D representation, we can easily project the multiview features into the structured 3D voxel space, which enables 3D native asset reconstruction. Compared with the native 3D representations such as voxel grid and triplane, the 2.5D representation is essentially images, and thus much easier to process. Hence, we can generate such 2.5D latents by adapting 2D multiview generative models as previous works [21, 40, 26].
@@
Figure 2: The multiview 2.5D representation. In this paper, we curate a 2.5D dataset by rendering the multiview RGB, normal and coordinate maps from 3D assets. With the dataset, we encode the image of each modalities into 2D latents with 2D VAE, and project the latents into 3D voxel space with multiview averaging. Then, we train our 3D refiner and decoder to reconstruct the 3D assets (3D gaussian splats or mesh) from the aggregated structured latents.
2 Related Works
2.1 2D Diffusion Models
@@ 2 Related Works
Diffusion models [35, 11] define a Markov process that gradually transforms data into noise through a forward diffusion process and then learns to reverse this transformation to generate data from noise. They have demonstrated remarkable performance in image [31, 9, 30, 32, 8] and video generation [12, 3, 4, 13]. However, conventional diffusion trajectories often exhibit significant curvature [17, 23], which increases the number of required solver steps and amplifies error accumulation. To address this, Rectified Flow [23, 1, 22] has been proposed, enforcing a straight-line trajectory between noise and data. Flow-based diffusion models have shown strong empirical performance and high efficiency, as demonstrated by recent state-of-the-art systems such as Stable Diffusion V3[7] and Flux.1 [2]. In this work, we adopt Flux.1 as the backbone diffusion model for 2.5D generation, leveraging its rectified flow design to enable high-quality and efficient image synthesis, which serves as a strong prior in our 3D reconstruction pipeline.
2.2 3D Generative Models
2.2.1 Multiview Diffusion Models
@@
Multiview diffusion models [47, 27, 28, 29, 5, 21, 34, 44] have focused on synthesizing geometrically consistent object views from varying camera poses, conditioned on image or text prompts. Early efforts in multiview generation models, such as MVDream [34], have successfully adapted pre-trained text-to-image diffusion models to generate object-centric multiview images. To enhance the geometric fidelity of the generated 3D content, methods such as Direct2.5 [27] and Kiss3DGen [21] incorporate geometry-aware designs and normal-guided supervision, producing multiview normal maps that facilitate high-quality 3D asset reconstruction. Other approaches leverage explicit 3D representations to improve realism and structural integrity.  Miao et al. [28], Qin et al. [29] distill multiview diffusion outputs into compact 3D generators using Gaussian splatting [18], enabling efficient single view 3D reconstruction with enhanced visual quality. In addition, auxiliary modules have been proposed to improve cross-view consistency and 3D reasoning.  Chen et al. [5] introduces a modular architecture called 3D-Adapter that injects 3D priors into diffusion pipelines, while  Zheng et al. [47] constructs a volumetric feature representation by aggregating multiview image features, leading to more accurate geometry reconstruction.
2.2.2 3D Diffusion Models
@@
Xiang et al. [41] introduces a unified structured latent space for 3D content (SLAT), supporting diverse representations like meshes, Gaussians, and NeRF.  Zhang et al. [46] uses a DiT-based latent diffusion model to generate neural surface representations and materials. It combines a multi-resolution VAE with PBR-aware training and supports conditional generation from text or images.  Li et al. [19] performs native 3D diffusion directly in volumetric space to generate coarse mesh shapes. It integrates a normal-guided surface refiner that enhances geometry.  Liu et al. [25] directly applies score-based diffusion to mesh vertex coordinates and face connectivity. It models 3D meshes in a native representation, enabling faithful and topology-preserving mesh generation without intermediate voxelization or rendering supervision.
Figure 3: Our proposed framework. In this paper, we propose a unified framework for (b) text or image to 3D generation and (c) geometry-conditioned texture generation. To provide the (a) hybrid image-text condition for 3D or texture generation, we adopt the off-the-shelf image encoder SigLIP [45] and Flux.1-Redux [2] image embedder with the T5 and CLIP text encoder. During the training, we randomly dropout the image or text condition to maintain the model’s ability to generate coherent content with ”X“ conditions and achieve better model performance, please refer to our experiment section for more details.
3 Method
@@
3.1 Bridging 3D and 2D with 2.5D Representation
In our framework, we utilize an image-like 2.5D representation (Fig. 2, stage I) for object encoding and decoding. Unlike 3D representations, the 2.5D representations are essentially images, however, it combines the multiview RGB, normal and coordinate images to represent the volume, surface and texture information that are sufficient for high-quality 3D objects reconstruction. Although some previous works [21, 40, 26, 38] have explored the potential of similar representations that composed of rgb with either coordinate or normal map, the absence of either normal or coordinate map in their framework could introduce extra difficulties in the 3D reconstruction stage.
From 2.5D to 3D. While most previous works adapted sparse-view reconstruction approaches to reconstruct the object from either rgb, normal or coordinate information, their result often limited to specific 3D representation, and sometimes low-quality in the unseen areas of the multi-view results. Inspired by the 3D structured latent representation proposed by Trellis [41], we also create a structure latent representation, which can then be decoded to various 3D representations in high-quality. For pipeline simplicity, we choose the VAE of the 2D diffusion model (i.e. Flux.1-dev) to encode the 2.5D images into 2D latents. Then, we project the 2D latents with $C$ channels and $N$ views $x_{2d}\in R^{N\times 3\times C\times H\times W}$ into a 3D latent voxel grid $x_{3d}\in R^{3C\times X\times Y\times Z}$ accordingly. As shown in stage II of Fig. 2, the projection of the 2D latent introduces a voxel-like 3D latent representation that contains the color, surface and coordinate information. Following [41], we aim to introduce a sparse 3D decoder to reconstruct the structured latent into explicit 3D representations such as 3D gaussian splats or mesh.
@@
$\hat{x}_{occ}=\begin{cases}1,&\text{if}~{}x_{occ}+b_{occ}>0.5\\ 0.&\text{otherwise}\end{cases}$ (1)
Then, we train the decoder to convert the refined structure latent into target 3D representation such as 3D gaussian splatting or mesh. Following  [41], we adopt the sparse transformer for decoder, and train the refiner unet and 3DGS decoder in an end-to-end manner from scratch with rendering loss only. For the mesh decoding, we fixed the pretrained refiner and train the mesh decoder only. Please refer to our appendix for more details.
3.2 Model Fine-tuning Via Mixture-of-LoRA
Inspired by Lin et al. [21], we aims to finetune a pretrained 2D diffusion model, i.e. Flux.1-dev [2], via Low-Rank Adaptation (LoRA) [15] for 2.5D image generation. As our 2.5D image comprise RGB, normal, and coordinate images, extending the 2D diffusion model to the joint generation of multi-modal images using LoRA adapters with limited data presents a challenge. Furthermore, we observe that simply finetune a single LoRA adapter with large rank (e.g. 384) do not achieves the best performance, both quantitatively and qualitatively (see Sec. 9).
Figure 4: The architecture of proposed modules. In order to fix the occluded areas in the structure latent aggregated from our 2.5D latent, we introduce the (a) 3D Residual UNet to refine the feature and occupancy field. Then, we apply the (b) sparse transformer decoder [41] for 3DGS or mesh reconstruction. (c) For each MLP block, we introduce three separate LoRA adapters tailored for 2.5D latent generation. Specifically, while the general LoRA layer is shared by all modalities, we feed the normal and coordinate (coord.) features into the auxiliary LoRA layers for extra feature projection, then we merge the results of the general and normal or coord. LoRAs via a simple addition.
@@ 3.2 Model Fine-tuning Via Mixture-of-LoRA
$h_{i}=\begin{cases}F(x_{i})+\hat{F}_{\text{general}}(x_{i}),&\text{if}~{}i=\text{``rgb''}\\ F(x_{i})+\hat{F}_{\text{general}}(x_{i})+\hat{F}_{i}(x_{i}).&\text{otherwise}\end{cases}$ (2)
@@
Joint training of text and image condition. Unlike previous works, we build a versatile framework for text and image conditioned generation of the 2.5D latent. As we chose the pretrained 2D diffusion model (i.e., Flux-dev.1) as the base model for 2.5D generation, we adopt the off-the-shelf text encoder (T5 and CLIP) and image encoder (SigLIP [45] image encoder with Flux-Redux [2] image embedder) to embed the text and image conditions respectively. During the model finetuning, we randomly drop the image or text condition, which we found beneficial to both text and image conditioned generation tasks. Please refer to our experiment section for more details.
4 Experiment
4.1 Training details
@@
4.2 Dataset
Due to the heterogeneous quality of the original Objaverse dataset, we initially excluded assets that lacked texture maps or exhibited low mesh fidelity. A comprehensive manual curation process was then performed to remove low-quality samples, including incomplete geometries, scanned planar surfaces, and large-scale scene compositions. Additionally, to address inconsistencies in object orientations, we manually annotated the canonical front-facing direction for each object. Then, we further apply a filtering following Trellis [41] to remove the items with low aesthetic score.
@@ 4.2 Dataset
For data rendering, we used Blender with a fixed camera distance of 4.5 units and a field of view (FoV) of 30 degrees. Each object was rendered from six distinct views to construct our 2.5D representation: four views were captured at 90-degree intervals in azimuth with a fixed elevation of 5 degrees—starting from the annotated front-facing direction—and an additional top-down view was rendered from a 90-degree elevation aligned with the frontal axis. For each view, we rendered RGB images, surface normal maps, and coordinate maps at a resolution of 512×512 pixels. The RGB multi-view images were then used as inputs for generating 3D-aware textual annotations via GPT-4V and Florence2 [42].
4.3 Evaluation
For evaluation, we conducted both quantitative and qualitative comparisons. Quantitatively, we used the Google Scanned Objects (GSO) dataset [6]. For the text-to-3D and text-to-texture task, we randomly sampled 100 objects and generated four-view renderings with GPT-4V annotations as text prompts. For image-to-3D and image-to-texture, we sampled 200 objects and rendered a single front-facing view as input for evaluation.
@@ 4.3 Evaluation
4.3.1 Metrics
We employ a comprehensive set of quantitative metrics to evaluate both the visual fidelity and geometric accuracy of the generated 3D models. For the text-to-3D and text-to-texture task, where precise 3D ground truth is unavailable, we assess semantic alignment using the CLIP score [10] by rendering four orthogonal views of each generated object. In addition, we use Q-Align [39], a large multi-modal evaluation model, to assess the quality and aesthetic consistency of the rendered outputs.
@@ 4.3.1 Metrics
4.4 Comparison with State-of-the-Art Methods
In this section, we compare our approach with several state-of-the-art methods across three tasks: text-to-3D generation, image-to-3D generation and texture generation. For text-to-3D generation, we compared our method with four representative baselines: the diffusion-based model CraftsMan [19] and Trellis [41], Hunyuan3D-1 [36] which builds on a large reconstruction model (LRM), Unique3D [40], a two stage generation method. For image-to-3D generation, We compared our method against three baselines: the latent diffusion model 3DTopia [14], Hunyuan3D-1 [36], and Trellis [41]. For texture generation, we compared our method against Hunyuan2.0 [37], MV-Adapter [16] which generates a single image from text, and then expand it into six multi-view images in image-to-texture generation, and SyncMVD [24] which generates a central image from text and predicts synchronized multi-view images via a cross-view fusion module, TexGaussian [43] using 3D Gaussians in text-to-texture generation.
Image-to-3D generation. As shown in Table 1, our method achieves competitive performance across both geometric and perceptual metrics. Notably, our results are comparable to Trellis, despite Trellis relying on significantly larger-scale training data and a much larger model. In contrast, our method is designed to be more data- and compute-efficient, yet achieves the highest PSNR and a competitive LPIPS score, ranking closely behind Trellis. Our SSIM score is also comparable to top-performing methods, indicating strong perceptual consistency and visual fidelity in the rendered views. These results suggest that our approach effectively reconstructs high-quality 3D shapes from single-view inputs, achieving a favorable trade-off between visual quality and computational cost. In addition to the quantitative results, Figure 5 presents a qualitative comparison under the image-to-3D setting. Compared to TripoSG [20] and Trellis, our method produces significantly more faithful and detailed textures. For example, our results better preserve color contrast, facial features, and fine-grained patterns such as fur, eyes, and accessories. Furthermore, our textures remain consistent across multiple viewpoints, reflecting better integration between appearance and structure.